A true-false test has 5 questions. What is the probability of getting at least 4 wrong by guessing all the answers?
Answer #1

probability of getting at least 4 wrong by guessing all the answers

=P(4 wrong)+P(5 wrong) =(5C4)*(1/5)4*(4/5)+(5C5)*(1/5)5*(4/5)0=20/3125+1/3125 =21/3125